OpenStudy (anonymous):

A regular box of cereal holds 14 ounces. If an economy box is made by doubling all the dimensions of the regular-sized box, how many ounces will the economy box hold?

OpenStudy (anonymous):

8 times bigger, 112 ounces

OpenStudy (anonymous):

the area of a box is x,y,z multiplied together, if you double those, 2x,2y,2z, it's a factor of 8 times larger.

OpenStudy (anonymous):

if the dimensions of the original box are L, B and H then volume = L*B*H = LBH = 14 oz ------- (1) on doubling the dimensions we have 2L, 2B, 2H so volume = 2L*2B*2H = 8 LBH = 8*14 {using (1)} = 112 oz